Find the equivalent fraction of 3 / 5 having
 denominator 30 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We require denominator 30
Let D be the numerator of the fraction
∴D / 30 = 3 / 5
5 × D = 3 × 30
D = (3 × 30) / 5
= 18
Therefore the required fraction is 18 / 30

Thanks